OpenStudy (anonymous):

DANGER! need help! please? please help! crashing for finals! find cos(u+v) given that sin u=3/5 and cos v=-5/13 (both u and v are in Quad 2

jimthompson5910 (jim_thompson5910):

sin u = 3/5 sin^2 u = 9/25 1 - sin^2 u = 1 - 9/25 cos^2 u = 16/25 cos u = -4/5 ... Note: we're in quad II -------------------------------- cos v = -5/13 cos^2 v = 25/169 1 - cos^2 v = 1 - 25/169 sin^2 v = 144/169 sin v = 12/13 ------------------------------------ cos(u+v) cos(u)cos(v)-sin(u)sin(v) (-4/5)(-5/13)-(3/5)(12/13) 4/13-36/65 -16/65 So cos(u+v) = -16/65

OpenStudy (anonymous):

i understand the first chunk that u typed but i got lost on that same part where u typed 1-sin^2 u=1-9/25 why did this have to be done?

jimthompson5910 (jim_thompson5910):

I want to go from sin u to cos u So I'm using the idea that sin^2 + cos^2 = 1 or 1 - sin^2 = cos^2

OpenStudy (anonymous):

ahhhh okay but why is it necessary to change to cosine?

jimthompson5910 (jim_thompson5910):

because cos(u+v) expands out to cos(u)cos(v)-sin(u)sin(v) so I need to find cos(u) using the given sin(u)

jimthompson5910 (jim_thompson5910):

I also needed to find sin(v) using the given cos(v)

OpenStudy (anonymous):

hmm okay i c then that means that we have to get the sin(v) using the cos right?

jimthompson5910 (jim_thompson5910):

exctly

jimthompson5910 (jim_thompson5910):

exactly*

OpenStudy (anonymous):

okay so i basically have to get the cos and sin given and get their opposites right and once i do that i add them>?

jimthompson5910 (jim_thompson5910):

you multiply them out like I wrote above, then you subtract

OpenStudy (anonymous):

o okay the original given cos and sin by the ones that were gathered right?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

YEAH! ALRIGHT I THINK I GOT THIS.. HEY IF U DON'T MIND....CAN U GIVE ME A COUPLE OF PROBLEMS TO PRACTICE ON MY OWN?

jimthompson5910 (jim_thompson5910):

find cos(u+v) given sin(u) = 3/5 and cos(v) = 8/17 (both u and v are in Quad I)

OpenStudy (anonymous):

OKAY THANK U I'LL USE IT TO PRACTICE LATER I APPRECIATE IT!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

:)